OpenStudy (anonymous):

The temperature of an oven went from 360° to 280°. What was the percent decrease in temperature to the nearest percent? A. 22% B. 29% C. 80% D. 130%

OpenStudy (anonymous):

a

OpenStudy (anonymous):

@AlexandervonHumboldt2 @Michele_Laino

OpenStudy (anonymous):

a

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (anonymous):

@AlexandervonHumboldt2

OpenStudy (anonymous):

@alivejeremy

OpenStudy (solomonzelman):

Example: Sam had 500 apples, but he ate some apples, and now Sam has only 350 apples left. Solution: \(\color{#000000 }{ \displaystyle 500-350=150 }\) (that is the amount of apples Sam ate, or alternatively it is the amount by which the number decreased) \(\color{#000000 }{ \displaystyle {\rm decrease~proportion}=\frac{500-150}{500}=\frac{35}{50} =0.7 }\) (Proportion of decrease in apples to the initial amount of apples) \(\color{#000000 }{ \displaystyle {\text %}~{\rm decrease~proportion}=0.7 \times 100=70{\text{%}} }\)

OpenStudy (anonymous):

wait

OpenStudy (alexandervonhumboldt2):

aweeee

OpenStudy (anonymous):

.2222222222222222222222222222222222222

OpenStudy (alexandervonhumboldt2):

no i made a msitake

OpenStudy (anonymous):

so yah i am correct thxs